Solve for t:
3 divided by t+3 minus 1 divided by t equals 5 divided by 2t plus 6
or
3/t+3-1/t=5/2t+6

Found 2 solutions by Alan3354, JoelSchwartz:
Answer by Alan3354(69443)


3/t+3-1/t=5/2t+6
----
6/(2t+6) - 1/t = 5/(2t+6)
1/(2t+6) = 1/t
2t+6 = t
t = -6

Answer by JoelSchwartz(130)

3/(t+3)-1/t=5/(2t+6)
(t+3)(3/(t+3)-1/t=5/(2t+6))
3-(t+3)/t=5(t+3)/(2t+6)
3-(t+3)/t=5(t+3)/2(t+3)
3-(t+3)/t=5/2
t(3-(t+3)/t=2.5)
3t-t-3=2.5t
2t-3=2.5t
-.5t=3
-t=6
t=-6